OpenStudy (anonymous):

4 4 ------ + ------ add b+3 b^2-9

OpenStudy (anonymous):

to add fractions, you need the same denominators in this case, you better factor that second denominator first then multiply by whatever it takes to make the denominators the same and then it is simple addition across the numerators

OpenStudy (anonymous):

i dont need to solve for b i need to add the two

OpenStudy (anonymous):

and for same denomenator it is b-3, b+3

OpenStudy (anonymous):

yes

OpenStudy (zepp):

Little hint there: (b+3)*(b-3) = b^2 - 3^2

OpenStudy (zepp):

Which is b^2 - 9

OpenStudy (anonymous):

4(b-3) 4 ------ + -------- (b+3)(b-3) (b+3)(b-3)

OpenStudy (zepp):

\[\frac{4(b-3)+4}{b^2 - 9}=\frac{4b-12+4}{b^2 - 9}=\frac{4b-8}{b^2 - 9}\]

OpenStudy (anonymous):

my possible asnwers are

OpenStudy (zepp):

\[\frac{4b-8}{b^2-9} = \frac{4(b-2)}{(b+3)(b-3)}\] :)